# Latent Geometry of Taste: Scalable Low-Rank Matrix Factorization

### v1

@@ 2.1 Exploratory Data Analysis
We further examine the distribution of explicit feedback values. As shown in Figure 2(a), the ratings are left-skewed, with a global mean $\mu\approx 3.53$. This indicates a positivity bias, where users are more inclined to rate items they enjoyed. Figure 2(c) highlights the exponential growth in content production over the decades, while Figure 2(d) demonstrates that the average rating has remained relatively stable despite this influx.
We analyze the semantic content of the dataset via genre distribution. Figure 2(f) highlights the class imbalance, with Drama and Comedy constituting the majority of the corpus. This imbalance poses a challenge for the model to learn distinct latent features for niche genres without overfitting to dominant categories.
@@
(b) Count of movie ratings per year.
(c) Movies released per year.
(d) Average movie rating over time.
(e) Distribution of genre showing outliers.
(f) Distribution of top movie genres.
@@ 2.2 Data Preprocessing
To facilitate matrix operations, we implemented an indexing layer that maps raw IDs to contiguous integers $u\in[0,N_{users}-1]$ and $i\in[0,M_{movies}-1]$. We employed a stratified train-test split strategy to preserve the user activity distribution. For every user, we randomly sample 80% of their ratings for training and hold out the remaining 20% for testing. This ensures that all users in the test set have a representation in the training phase, mitigating the user cold-start problem during evaluation.
@@ 3.5 Hyperparameter Tuning
We conduct an exhaustive Grid Search to identify the optimal configuration for training. We explore the space shown in Table 1. We trained separate models for every combination of $k,\lambda,\tau$ for 20 epochs. To maintain computational feasibility during the search, Precision and Recall metrics are evaluated on a random sample of 3,000 users, while RMSE is calculated on the full test set. We then analyze the trade-off between model complexity and generalization performance to select the final model parameters.
Table 1: Hyperparameter search space
@@ 4 Results and Conclusion
Figure 4: Training dynamics of the best performing model.
Table 2: Hyperparameter grid search results
@@ 4.1 Movie Trait Vectors
Figure 5: Visualization of latent trait vectors for selected movies.
@@ Appendix A Movie Recommender Application
The best performing model was deployed as a Streamlit application. Figure 6 illustrates the application interface, where users can search the full movie catalog, assign ratings to movies they have watched, and receive personalized recommendations inferred from their expressed preferences.

### v2

@@ 2.1 Exploratory Data Analysis
We further examine the distribution of explicit feedback in the MovieLens 32M dataset. As illustrated in Figure 2(a), the ratings exhibit a left-skewed distribution, with a global mean of approximately 3.53. This pattern reflects a positivity bias, whereby users tend to assign ratings primarily to movies they appreciated. Figures 2(c) and 2(b) reveal exponential growth in movie releases and user ratings, respectively, across decades, followed by a sharp decline in recent years, potentially attributable to dataset collection artifacts or shifts in user behavior. Despite these trends, Figure 2(d) indicates that average ratings have remained relatively stable over time, fluctuating modestly around the global mean.
To explore semantic patterns, we assess genre related characteristics. Figure 2(e) depicts the average ratings across top genres, revealing variations such as higher means for Drama and Documentary compared to Horror. Figure 3 extends this analysis by plotting average ratings against genre popularity, highlighting that niche genres like Film-Noir achieve high ratings despite lower popularity, while mainstream genres like Comedy show moderate ratings with high engagement. Figure 2(f) underscores genre imbalance, with Drama (20.5%) and Comedy (13.9%) dominating the corpus, followed by Thriller (7.1%), Romance (6.2%), and others; this skew challenges models to capture nuanced features for underrepresented categories without bias toward dominant ones. Finally, Figure 4 presents a word cloud of user-assigned tags, emphasizing prevalent themes such as atmospheric, sci-fi, comedy, and action, which provide additional semantic context for movie representations.
@@
(b) Distribution of movie rating values.
(c) Number of movies released per year.
(d) Average movie rating by release year.
(e) Average ratings for top movie genres.
(f) Percentage distribution of top movie genres.
@@
Figure 3: Average rating versus popularity for movie genres
@@ 2.2 Data Preprocessing
To facilitate matrix operations, we implemented an indexing layer that maps raw identifiers to contiguous integers $u\in[0,N_{users}-1]$ and $i\in[0,M_{movies}-1]$. We employed a stratified train-test split strategy to preserve the user activity distribution. For every user, we randomly sample 80% of their ratings for training and hold out the remaining 20% for testing. This ensures that all users in the test set have a representation in the training phase, mitigating the user cold-start problem during evaluation.
@@ 3.5 Hyperparameter Tuning
We conduct an exhaustive Grid Search to identify the optimal configuration for training. We explore the space shown in Table 1. We trained separate models for every combination of $k,\lambda,\tau$ for 20 epochs. To maintain computational feasibility during the search, Precision and Recall metrics are evaluated on a random sample of 3,000 users, while RMSE is calculated on the full test set. We then analyze the trade-off between model complexity and generalization performance to select the final model parameters. All models were trained using complementary compute provided on Google Colab.
Table 1: Hyperparameter search space explored for model training
@@ 4 Results and Conclusion
Figure 5: Training dynamics of the best performing model (k=10, $\lambda$=0.1, $\tau$=0.25).
Table 2: Hyperparameter grid search results for the MF model, including Training RMSE and Test RMSE
@@ 4.1 Movie Trait Vectors
Figure 6: Visualization of latent trait vectors for selected movies, projected into 2D space using PCA.
@@ Appendix A Movie Recommender Application
The best performing model was deployed as a Streamlit application. Figure 7 illustrates the application interface, where users can search the full movie catalog, assign ratings to movies they have watched, and receive personalized recommendations inferred from their expressed preferences. Live demo can be found here: https://als-recommender.streamlit.app/